I NEED HELP IN THIS QUESTION ASAP PLEASEA Transmitter provides 15 W to an antenna having 12 dB gain. The receiver Antenna has a gain of 3 dB and the receiver bandwid
Answer #1

NF=8 dB, BW= 30 KHZ Noise floor (dBm) = 10 log (KX To X1000) + NF t . - 10109,0 BW = 10 logio (1038 X10-23 x 290 x1000) +8 +then log (2) = -1.019 - 101-89 = 1.508 + 40 Now , 0.95 - 012) = 1 - Q(-2) = Q(-) = 0.05 o, -d = 1.65 1. Q = -1.65 és log = 1.